Three farmers (Farms A, B, and C) have offered their land for sale for business development....

  1. Three farmers (Farms A, B, and C) have offered their land for sale for business development. Each parcel of land is approximately 1,000 acres and all the farms are located within the same region. The economic development committee in your small town has asked you to do some research on the farmer’s land before negotiations begin so the committee will know what to expect in advance as far as offering fair compensation. The farmers have made last year’s financial records available for their farms (they all grew soybeans) and your job is to assess the fertility of each farmer’s land. Assume each farmer uses current best practices regarding input costs.
  1. Use the farm records below to rank the farmers’ land with either a “top”, “medium” or “low” compensation ranking so that each farm receives a ranking. Note that this approach will follow the Ricardo theory of land value. (15 Points)

Farm A:

Average market price for the crops: $10.50/bushel

Average Yield: 34 bushel/acre

Average Production Cost per Acre: $162

Rank:

Farm B:

Average market price for the crops: $10.50/bushel

Average Yield: 38 bushel/acre

Average Production Cost per Acre: $132

Rank:

Farm C:

Average market price for the crops: $10.50/bushel

Average Yield: 41 bushel/acre

Average Production Cost per Acre: $157

Rank:

  1. Do you think that fertility alone should explain the value of land or are there other factors at play that should go into the compensation offer? Why or why not? (10 points)
Answer #1

Answer)

A) According to Ricardo, land's quality is owed to its original and indestructible properties and that its fertility is derived on the basis of diminishing marginal returns, keeping total cost pf production for each land same.

Owing to this gross idea we have, for all three lands, the average cost of production per acre and the average yield per acre. The return on the cost will give us the idea of fertility for each (since each farmer uses current best practices, so the effect of external factors could be reliably assumed cpnstant thus making the yield-cost ratio an indicative of fertility).

Land A: Bushel per dollar= 34/162= 0.21 bu/$ (approx.)

Land B: Bushel per dollar= 38/132= 0.29 bu/$ (approx.)

Land C: Bushel per dollar= 41/157= 0.26 bu/$ (approx.)

So Land B is top rank land,

Land C is medium rank land and,

Lsnd A is low rank land.

B) No, fertility alone cannot explain value of land. This was also one of the arguments of criticism of Ricardo's theory. A land is not only used for cultivation but is also used for constructing buildings, production plants, etc. and thus further used for commercial purposes. In such cases, the purpose of land changes and so fertility is no appropriate standard to assess its worth. Then, the land's location, its proximity to market, water sources etc, and even the dimensions contribute in deciding its value.